OpenStudy (anonymous):

How do I complete this fraction problem? 3/4 + 5/6

OpenStudy (anonymous):

what do you need? their sum?

OpenStudy (anonymous):

find common denominator. what can 4 and 6 both go into evenly?

OpenStudy (anonymous):

@Evern: I need the sum. @ Moldy84: 2

OpenStudy (anonymous):

make the denominator smae by multiplying numer and denom by a particular number

OpenStudy (anonymous):

that's the lowest common factor mathhelp. we are looking for a number that both of them can divide into. 12 is the commond denominator

OpenStudy (anonymous):

LCM of denominators is 12 so we have 9/12 + 10/12 = 19/12 which is 1 7/12

OpenStudy (anonymous):

3*6/24 +4*5/24 = 38/24

OpenStudy (anonymous):

so if we make both denomiators 12 we need to multiply the numerators by how much we multiplied the denominator on that particular fraction

OpenStudy (anonymous):

so (3*3)/12+(5*2)/12=(9/12+10/12)=19/12

OpenStudy (anonymous):

Thanks jimmyrep, Evern, siddharth tiwari, and moldy84!

OpenStudy (anonymous):

just gotta make sure to multiply the numerator by the same amount you multiplied the denominator or you are changing the fraction

OpenStudy (anonymous):

welcome